# Mastering Multi-Drone Volleyball through Hierarchical Co-Self-Play Reinforcement Learning

### v4

@@ Abstract
In this paper, we tackle the problem of learning to play 3v3 multi-drone volleyball, a new embodied competitive task that requires both high-level strategic coordination and low-level agile control. The task is turn-based, multi-agent, and physically grounded, posing significant challenges due to its long-horizon dependencies, tight inter-agent coupling, and the underactuated dynamics of quadrotors. To address this, we propose Hierarchical Co-Self-Play (HCSP), a hierarchical reinforcement learning framework that separates centralized high-level strategic decision-making from decentralized low-level motion control. We design a three-stage population-based training pipeline to enable both strategy and skill to emerge from scratch without expert demonstrations: (I) training diverse low-level skills, (II) learning high-level strategy via self-play with fixed low-level skills, and (III) joint fine-tuning through co-self-play. Experiments show that HCSP achieves superior performance, outperforming non-hierarchical self-play and rule-based hierarchical baselines with an average 82.9% win rate and a 71.5% win rate against the two-stage variant. Moreover, co-self-play leads to emergent team behaviors such as role switching and coordinated formations, demonstrating the effectiveness of our hierarchical design and training scheme. The project page is at https://sites.google.com/view/hi-co-self-play.
@@ 1 Introduction
To address these challenges, we propose Hierarchical Co-Self-Play (HCSP), a hierarchical reinforcement learning (HRL) framework that separates strategic reasoning from motor control. The policy is decomposed into a high-level strategy responsible for team tactics and multiple low-level skills responsible for drone-specific motion control. The high-level strategy is a centralized multi-layer perceptron (MLP) with three output heads, each producing commands for one drone, enabling synchronized decision-making across the team. The low-level policies are independent across drones and operate at a higher frequency to ensure responsive, fine-grained control. To reflect the turn-based structure of the game, the high-level strategy is only activated upon discrete game events–when the ball is struck or crosses the net–while low-level control runs continuously between these moments.
@@ 4.1 Game Results Against Baselines
Figure 4: Experiment results. (a) HCSP performance against baseline policies. (b) Cross-play win rate heatmap among ten randomly picked policies. (c) Ablation study on policy chaining in Stage I: policy chaining enables a successful transition from Attack to Hover, whereas single-policy fails.
@@ 6 Conclusion
Our approach is currently limited to high-fidelity simulation, which cannot fully capture real-world sensor noise or hardware latency. Consequently, HCSP-trained policies may encounter a sim-to-real gap and are unlikely to transfer zero-shot to physical drones. Nonetheless, we report preliminary sim-to-real results in Appendix G, which provide promising initial validation in real-world experiments. Furthermore, HCSP assumes access to full state information (drone poses, velocities, and ball dynamics) rather than raw sensor or vision inputs, restricting its applicability in settings where only onboard cameras are available.
@@ Acknowledgments
This research was supported by National Natural Science Foundation of China (No.62406159, 62325405), Postdoctoral Fellowship Program of CPSF under Grant Number (GZC20240830, 2024M761676), China Postdoctoral Science Special Foundation 2024T170496. Additional support was provided by Tsinghua-Efort Joint Research Center for EAI Computation and Perception, Beijing National Research Center for Information Science and Technology (BNRist), Beijing Innovation Center for Future Chips, and State Key laboratory of Space Network and Communications.
@@ A.2 Drone
where $\bm{x}_{W}$ and $\bm{v}_{W}$ are the position and velocity of the drone in the world frame, $\bm{R}_{WB}$ is the rotation matrix from the body frame to the world frame, $\bm{J}$ is the diagonal inertia matrix, $\bm{g}$ represents gravity, $\bm{q}$ is attitude in quaternion, and $\bm{\omega}$ is the angular velocity. The operator for quaternion multiplication is denoted by $\otimes$. External forces $\bm{F}$ include aerodynamic drag and downwash effects. The collective thrust $\bm{f}$ and torque $\bm{\eta}$ are calculated based on the thrust of each rotor $\bm{f}_{i}$ as:
@@ B.1.1 Serve
Observation and reward. The drone’s observation is a vector of dimension $37$ including the drone’s root state, the ball’s position, the ball’s relative position to the drone, the ball’s linear velocity, a one‑hot flag indicating whether the drone is permitted to hit the ball, and the ball’s relative position to the target ball position. The detailed specification of the task’s reward function per time step can be found in Table 3.
@@
Before
Hit
dist_to_ball $[0,0.16]$ drone’s distance to the ball
@@
After Hit
dist_to_anchor $[0,16]$ ball’s distance to anchor
End
Reward
in_side $\{0,10\}$ ball landing in opponent’s court
@@ B.1.2 Receive
Observation and reward. The drone’s observation is a vector of dimension $34$ including the drone’s root state, the ball’s position, the ball’s relative position to the drone, the ball’s linear velocity, and a one‑hot flag indicating whether the drone is permitted to hit the ball. The detailed specification of the task’s reward function per time step can be found in Table 4.
@@
Before
Hit
dist_to_ball $[0,10]$ drone’s distance to the ball
@@
After
Hit
dist_to_anchor $[0,50]$ ball’s distance to anchor
@@
End
Reward
in_side $\{0,10\}$ ball landing in opponent’s court
@@ B.1.3 Set
Observation and reward. The drone’s observation is a vector of dimension $34$ including the drone’s root state, the ball’s position, the ball’s relative position to the drone, the ball’s linear velocity, and a one‑hot flag indicating whether the drone is permitted to hit the ball. The detailed specification of the task’s reward function per time step can be found in Table 5.
@@
Before
Hit
dist_to_ball $[0,2]$ drone’s distance to the ball
@@
After
Hit
dist_to_anchor $[0,10]$ ball’s distance to anchor
@@
End
Reward
highest_ball_pos $[0,400]$ ball’s peak height exceeding 3m
@@ B.1.4 Attack
Observation and reward. The drone’s observation is a vector of dimension $36$ including the drone’s root state, the ball’s position, the ball’s relative position to the drone, the ball’s linear velocity, a one‑hot flag indicating whether the drone is permitted to hit the ball, and a one‑hot flag indicating the attack direction. The detailed specification of the task’s reward function per time step can be found in Table 6.
@@
Before
Hit
dist_to_ball $[0,1]$ drone’s distance to the ball
@@
After
Hit
dist_to_anchor $[0,150]$ ball’s distance to anchor
@@
End
Reward
ball_hit_ground $\{0,5\}$ ball hit the ground successfully
@@ B.1.5 Pass
Observation and reward. The drone’s observation is a vector of dimension $36$ including the drone’s root state, the ball’s position, the ball’s relative position to the drone, the ball’s linear velocity, a one‑hot flag indicating whether the drone is permitted to hit the ball, and a one‑hot flag indicating the opponent’s attacking direction. The detailed specification of the task’s reward function per time step can be found in Table 7.
@@
Before
Hit
dist_to_ball $[0,10]$ drone’s distance to the ball
@@
After
Hit
dist_to_anchor $[0,10]$ ball’s distance to anchor
@@
End
Reward
highest_ball_pos $[0,200]$ ball’s peak height exceeding 3m
@@ B.1.6 Hover
Observation and reward. The drone’s observation is a vector of dimension $26$ including the drone’s relative position to the target point, the drone’s root state except position, and the drone’s relative orientation to the ideal state. The detailed specification of the task’s reward function per time step can be found in Table 8.
@@
Hover
Reward
drone_pos $[0,3]$ drone’s relative position to the target point
@@ B.1.7 GoTo
Observation and reward. The drone’s observation is a vector of dimension $26$ including the drone’s relative position to the target point, the drone’s root state except position, and the drone’s relative orientation to the ideal state. The detailed specification of the task’s reward function per time step can be found in Table 9.
@@
Hover
Reward
drone_pos $[0,1]$ drone’s relative position to the target point
@@ C.2 Training Result
The results of high‑level strategy training are shown in Fig. 10. In our experiments, we set the maximum population size to five, yielding five iterative training phases; each phase terminates when either a preset number of training steps is reached or a win‑rate threshold is met. The lower triangular portion of the matrix reports the win rate of each row policy against each column policy (i.e., the probability that the row policy prevails). Because the game is symmetric and zero‑sum, the upper‑triangular entries are complementary, such that each pair of win rates sums to one.
@@ D.1 Ablation Study on Low-Level Skill Refinement Schedule
0: Low-level skill set $\{\pi^{L}\}$, high-level policy $\pi^{H}$, number of max iterations $M$, checkpoint saving interval NN, win rate threshold $\tau$
@@
4:  ${\pi_{k}^{L}}^{\prime}\leftarrow\pi_{k}^{L}$
5:  ${\pi_{k}^{H}}^{\prime}\leftarrow\pi^{H}$
6:  Form team $(\{\pi^{L}\}\setminus\{\pi_{k}^{L}\})\cup\{\pi^{L^{\prime}}_{k}\}$ with high-level $\pi^{H^{\prime}}_{k}$
7:  for $i=1$ to $M$ do
8:   Jointly optimize $\pi_{k}^{L}$$^{\prime}$ and $\pi_{k}^{H}$$^{\prime}$ through self-play against the frozen team $\{\pi^{L}\}$ and $\pi^{H}$
9:   if $i\bmod N=0$ then
10:    Save a checkpoint of $\pi_{k}^{L}$$^{\prime}$ and $\pi_{k}^{H}$$^{\prime}$
11:   end if
12:  end for
13:  For each $\pi^{L^{\prime}}_{k},\pi^{H^{\prime}}_{k}$ checkpoint, form team $(\{\pi^{L}\}\setminus\{\pi_{k}^{L}\})\cup\{\pi^{L^{\prime}}_{k}\}$ with high-level $\pi^{H^{\prime}}_{k}$ and evaluate the win rate against the frozen team $\{\pi^{L}\}$ and $\pi^{H}$.
14:  Select the checkpoint of $\pi_{k}^{L}$$^{\prime}$ with the highest win rate, denoted as $\pi^{L}_{k,\text{best}}$
15:  if win rate of $\pi^{L}_{k,\text{best}}>\tau$ then
16:   Append $\pi^{L}_{k,\text{best}}$ to $\Pi_{\text{new}}$
17:  end if
18:end for
19:return $\{\pi^{L}\}\cup\Pi_{\text{new}}$
@@ D.3 Ablation Study on High-Level Strategy Parameter Initialization
As the previous high-level strategy becomes invalid due to changes in the size of the low-level skill pool, it is necessary to retrain the high-level strategy to accommodate the augmented skill set. The initialization of this new high-level strategy is a critical factor influencing the overall performance in Stage III. We compare two different initialization methods for the high-level strategy. The first is a straightforward approach that trains the high-level strategy from scratch. In the second approach, we expand each actor head’s parameter vector to match the larger action space by copying over its original values. Any newly added dimensions are then initialized to the mean of that head’s original parameters. In both settings, we train the new high-level strategy with frozen augmented low-level skills against the frozen high-level strategy and low-level skills obtained after Stage II. As shown in Fig. 11(c), although training from scratch requires more time steps to converge, it ultimately achieves a higher win rate. In contrast, the parameter-averaging initialization method tends to converge to a suboptimal local solution.
@@ Appendix E Baselines
To ensure fairness, we provide all baseline agents with the full-state observations of opponent drones as those available to HCSP. Additionally, HCSP is trained independently without exposure to any of the baseline agents, guaranteeing an unbiased evaluation.
@@ F.2 Nash-Averaging Evaluator
While head-to-head win rates fail to reflect absolute policy strength in non-transitive games, and Elo ratings [22] exhibit sensitivity to the composition of the policy population, the Nash-averaging evaluator, introduced by Liu et al. [21], provides a robust alternative. This method selects a randomly sampled set of policies, then constructs a payoff matrix from pairwise match-ups, and finally computes a Nash equilibrium (NE). The resulting equilibrium weights yield a comparatively population-invariant measure of relative performance, mitigating biases induced by cyclic dominance. So we adopt this as an evaluation metric. The details are as follows.

### v5

@@ Abstract
In this paper, we tackle the problem of learning to play 3v3 multi-drone volleyball, a new embodied competitive task that requires both high-level strategic coordination and low-level agile control. The task is turn-based, multi-agent, and physically grounded, posing significant challenges due to its long-horizon dependencies, tight inter-agent coupling, and the underactuated dynamics of quadrotors. To address this, we propose Hierarchical Co-Self-Play (HCSP), a hierarchical reinforcement learning framework that separates centralized high-level strategic decision-making from decentralized low-level motion control. We design a three-stage population-based training pipeline to enable both strategy and skill to emerge from scratch without expert demonstrations: (I) training diverse low-level skills, (II) learning high-level strategy via self-play with fixed low-level controllers, and (III) joint fine-tuning through co-self-play. Experiments show that HCSP achieves superior performance, outperforming non-hierarchical self-play and rule-based hierarchical baselines with an average 82.9% win rate and a 71.5% win rate against the two-stage variant. Moreover, co-self-play leads to emergent team behaviors such as role switching and coordinated formations, demonstrating the effectiveness of our hierarchical design and training scheme. The project page is at https://hi-co-self-play.github.io.
@@ 1 Introduction
To address these challenges, we propose Hierarchical Co-Self-Play (HCSP), a hierarchical reinforcement learning (HRL) framework that separates strategic reasoning from motor control. The policy is decomposed into a high-level strategy responsible for team tactics, and multiple low-level skills responsible for drone-specific motion control. The high-level strategy is a centralized multi-layer perceptron (MLP) with three output heads, each producing commands for one drone, enabling synchronized decision-making across the team. The low-level policies are independent across drones and operate at a higher frequency to ensure responsive, fine-grained control. To reflect the turn-based structure of the game, the high-level strategy is only activated upon discrete game events–when the ball is struck or crosses the net–while low-level control runs continuously between these moments.
@@ 4.1 Game Results Against Baselines
Figure 4: Experiment results. (a) HCSP performance against baseline policies. (b) Cross-play win rate heatmap among ten randomly picked policies. (c) Ablation study on policy chaining in Stage I: policy chaining enables successful transition from Attack to Hover, whereas single-policy fails.
@@ 6 Conclusion
Our approach is currently limited to high-fidelity simulation, which cannot fully capture real-world sensor noise, or hardware latency. Consequently, HCSP-trained policies may encounter a sim-to-real gap and are unlikely to transfer zero-shot to physical drones. Nonetheless, we report preliminary sim-to-real results in Appendix G, which provide promising initial validation in real-world experiments. Furthermore, HCSP assumes access to full state information (drone poses, velocities, and ball dynamics) rather than raw sensor or vision inputs, restricting its applicability in settings where only onboard cameras are available.
@@ Acknowledgments
This research was supported by National Natural Science Foundation of China (No.62406159, 62325405), Postdoctoral Fellowship Program of CPSF under Grant Number (GZC20240830, 2024M761676), China Postdoctoral Science Special Foundation 2024T170496.
@@ A.2 Drone
where $\bm{x}_{W}$ and $\bm{v}_{W}$ are the position and velocity of the drone in the world frame, $\bm{R}_{WB}$ is the rotation matrix from the body frame to the world frame, $\bm{J}$ is the diagonal inertia matrix, $\bm{g}$ represents gravity, $\bm{q}$ is attitude in quaternion, and $\bm{\omega}$ is the angular velocity. The operator for quaternion multiplication is denoted by $\otimes$. External forces $\bm{F}$ includes aerodynamic drag and downwash effects. The collective thrust $\bm{f}$ and torque $\bm{\eta}$ are calculated based on the thrust of each rotor $\bm{f}_{i}$ as:
@@ B.1.1 Serve
Observation and reward. The drone’s observation is a vector of dimension $37$ including the drone’s root state, the ball’s position, the ball’s relative position to the drone, the ball’s linear velocity, a one‑hot flag indicating whether the drone is permitted to hit the ball, the ball’s relative position to the target ball position. The detailed specification of the task’s reward function per time step can be found in Table 3.
@@
Before Hit dist_to_ball $[0,0.16]$ drone’s distance to the ball
@@
After Hit dist_to_anchor $[0,16]$ ball’s distance to anchor
End Reward in_side $\{0,10\}$ ball landing in opponent’s court
@@ B.1.2 Receive
Observation and reward. The drone’s observation is a vector of dimension $34$ including the drone’s root state, the ball’s position, the ball’s relative position to the drone, the ball’s linear velocity, a one‑hot flag indicating whether the drone is permitted to hit the ball. The detailed specification of the task’s reward function per time step can be found in Table 4.
@@
Before Hit dist_to_ball $[0,10]$ drone’s distance to the ball
@@
After Hit dist_to_anchor $[0,50]$ ball’s distance to anchor
@@
End Reward in_side $\{0,10\}$ ball landing in opponent’s court
@@ B.1.3 Set
Observation and reward. The drone’s observation is a vector of dimension $34$ including the drone’s root state, the ball’s position, the ball’s relative position to the drone, the ball’s linear velocity, a one‑hot flag indicating whether the drone is permitted to hit the ball. The detailed specification of the task’s reward function per time step can be found in Table 5.
@@
Before Hit dist_to_ball $[0,2]$ drone’s distance to the ball
@@
After Hit dist_to_anchor $[0,10]$ ball’s distance to anchor
@@
End Reward highest_ball_pos $[0,400]$ ball’s peak height exceeding 3m
@@ B.1.4 Attack
Observation and reward. The drone’s observation is a vector of dimension $36$ including the drone’s root state, the ball’s position, the ball’s relative position to the drone, the ball’s linear velocity, a one‑hot flag indicating whether the drone is permitted to hit the ball, a one‑hot flag indicating the attack direction. The detailed specification of the task’s reward function per time step can be found in Table 6.
@@
Before Hit dist_to_ball $[0,1]$ drone’s distance to the ball
@@
After Hit dist_to_anchor $[0,150]$ ball’s distance to anchor
@@
End Reward ball_hit_ground $\{0,5\}$ ball hit the ground successfully
@@ B.1.5 Pass
Observation and reward. The drone’s observation is a vector of dimension $36$ including the drone’s root state, the ball’s position, the ball’s relative position to the drone, the ball’s linear velocity, a one‑hot flag indicating whether the drone is permitted to hit the ball, a one‑hot flag indicating the opponent’s attacking direction. The detailed specification of the task’s reward function per time step can be found in Table 7.
@@
Before Hit dist_to_ball $[0,10]$ drone’s distance to the ball
@@
After Hit dist_to_anchor $[0,10]$ ball’s distance to anchor
@@
End Reward highest_ball_pos $[0,200]$ ball’s peak height exceeding 3m
@@ B.1.6 Hover
Observation and reward. The drone’s observation is a vector of dimension $26$ including the drone’s relative position to the target point, the drone’s root state except position, the drone’s relative orientation to the ideal state. The detailed specification of the task’s reward function per time step can be found in Table 8.
@@
Hover Reward drone_pos $[0,3]$ drone’s relative position to the target point
@@ B.1.7 GoTo
Observation and reward. The drone’s observation is a vector of dimension $26$ including the drone’s relative position to the target point, the drone’s root state except position, the drone’s relative orientation to the ideal state. The detailed specification of the task’s reward function per time step can be found in Table 9.
@@
Hover Reward drone_pos $[0,1]$ drone’s relative position to the target point
@@ C.2 Training Result
The results of high‑level strategy training are shown in Fig. 10. In our experiments, we set the maximum population size to five, yielding five iterative training phases; each phase terminates when either a preset number of training steps is reached or a win‑rate threshold is met. The lower triangular portion of the matrix reports the win rate of each row policy against each column policy (i.e. the probability that the row policy prevails). Because the game is symmetric and zero‑sum, the upper‑triangular entries are complementary, such that each pair of win rates sums to one.
@@ D.1 Ablation Study on Low-Level Skill Refinement Schedule
0: Low-level skill set $\{\pi^{L}\}$, high-level policy $\pi^{H}$, number of max iterations $M$, checkpoint saving interval NN, win rate threshold $t$
@@
4:  ${\pi^{L}}^{\prime}\leftarrow\pi^{L}$
5:  ${\pi^{H}}^{\prime}\leftarrow\pi^{H}$
6:  for $i=1$ to $M$ do
7:   Jointly optimize $\pi^{L}$$^{\prime}$ and $\pi^{H}$$^{\prime}$ through self-play against frozen $\{\pi^{L}\}$ and $\pi^{H}$
8:   if $i\bmod N=0$ then
9:    Save a checkpoint of $\pi^{L}$$^{\prime}$ and $\pi^{H}$$^{\prime}$
10:   end if
11:  end for
12:  Evaluate all saved checkpoints of $\pi^{L}$$^{\prime}$ and $\pi^{H}$$^{\prime}$ win rates against frozen $\{\pi^{L}\}$ and $\pi^{H}$
13:  Select the checkpoint of $\pi^{L}$$^{\prime}$ with the highest win rate, denoted as $\pi^{L}_{\text{best}}$
14:  if win rate of $\pi^{L}_{\text{best}}>t$ then
15:   Add $\pi^{L}_{\text{best}}$ to $\Pi_{\text{new}}$
16:  end if
17:end for
18:return $\{\pi^{L}\}\cup\Pi_{\text{new}}$
@@ D.3 Ablation Study on High-Level Strategy Parameter Initialization
As the previous high-level strategy becomes invalid due to changes in the size of the low-level skill pool, it is necessary to retrain the high-level strategy to accommodate the augmented skill set. The initialization of this new high-level strategy is a critical factor influencing the overall performance in Stage III. We compare two different initialization methods for the high-level strategy. The first is a straightforward approach that trains the high-level strategy from scratch. In the second approach, we expand each actor head’s parameter vector to match the larger action space by copying over its original values. Any newly added dimensions are then initialized to the mean of that head’s original parameters. In both settings, we train the new high-level strategy against the frozen high-level strategy and low-level skills obtained after Stage II. As shown in Fig. 11(c), although training from scratch requires more time steps to converge, it ultimately achieves a higher win rate. In contrast, the parameter-averaging initialization method tends to converge to a suboptimal local solution.
@@ F.2 Nash-Averaging Evaluator
While head-to-head win rates fail to reflect absolute policy strength in non-transitive games, and Elo ratings [6] exhibit sensitivity to the composition of the policy population, Nash-averaging evaluator, introduced by Liu et al. [14], provides a robust alternative. This method selects a randomly sampled set of policies, then constructs a payoff matrix from pairwise match-ups, and finally computes a Nash equilibrium (NE). The resulting equilibrium weights yield a comparatively population-invariant measure of relative performance, mitigating biases induced by cyclic dominance. So we adopt this as an evaluation metric. The details are as follows.